OpenStudy (anonymous):

does this work: r=-4 sin theta 4=-r sin theta Something makes me feel like this doesn't work but I have to check.

OpenStudy (anonymous):

what's your purpose?

OpenStudy (anonymous):

trying to convert from polar form to cartesian. My next step would be to substitute x in where x=r cos theta

OpenStudy (anonymous):

No. it doesn't work.

OpenStudy (amistre64):

to check, consider this: 30 = 6 * 5 6 = -30 * 5

OpenStudy (amistre64):

cartesian needs an r^2 and some sort of r sin, or r cos to play with

OpenStudy (amistre64):

(r=-4 sin(t) ) *r r^2 = -4 rsin(t) since r^2 = x^2 + y^2; and y = r sin(t) we get: x^2 + y^2 = -4y or so ive been led to believe